OpenStudy (anonymous):

A snow cone is a tasty treat with flavored ice and a spherical bubble gum ball at the bottom. Snow cone with spherical bubble gum ball at the bottom The radius of the cone is 1.5 inches, and its height is 3 inches. If the diameter of the bubble gum ball is 0.5 inches, what is the closest approximation of the volume of the cone that can be filled with flavored ice? 21.21 in3 0.07 in3 7.00 in3 7.07 in3

OpenStudy (anonymous):

@ganeshie8 @robtobey

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

So basically you want to subtract the volume of the sphere from the volume of the cone. The volume of a sphere is V=(4/3)*pi*r^3 ---- we're given d=0.5, and d=2r so r=0.5/2=0.25; use this to find V The volume of a cone is V=(1/3)*pi*r^2*h ---- we're given r=1.5 and h=3 so plug them in. Once you've found both V's subtract the first one from the second one to get your answer. so it's 7.00in^3

OpenStudy (anonymous):

Can I please ask another question if u don't mind :)

OpenStudy (anonymous):

sure just could i have the medal for this one?

OpenStudy (anonymous):

and put it in a seperate question

OpenStudy (anonymous):

OK thank you I have a couple questions ur very helpful and im very thankful